I am having difficulty solving this problem! I know it has already been answered, but the values of the equation that has been used is different than mine. PLEASE HELP!!!
Suppose you are an event coordinator for a large performance theater. One of the hottest new Broadway musicals has started to tour and your city is the first stop on the tour. You need to supply information about projected ticket sales to the box office manager. The box office manager uses this information to anticipate staffing needs until the tickets sell out.
You provide the manager with a quadratic equation that models the expected number of ticket sales for each day x.
(x = 1 is the day tickets go on sale).
Tickets = -0.6x2 + 12x + 11
Does the graph of the equation above open up or down?
How did you determine whether the graph of the equation above opened up or down?
Describe what happens to the ticket sales as time passes.
Use the quadratic formula to determine the last day that tickets will be sold.
Hint: Write your answer in terms of the number of days after ticket sales begin and you can’t have partial days.
Will tickets peak or be at a low during the middle of the sale?
How do you know?
After how many days will the peak or low occur?
Hint: Use the axis of symmetry formula.
How many tickets will be sold on the day when the peak or low occurs?
Hint: Substitute your value from question 7 into the original quadratic equation.
What are the coordinates of the vertex?
How did you determine this?
How many solutions are there to the original quadratic equation given in question 3?
How do you know?




Answer by nerdybill(7384)

Suppose you are an event coordinator for a large performance theater. One of the hottest new Broadway musicals has started to tour and your city is the first stop on the tour. You need to supply information about projected ticket sales to the box office manager. The box office manager uses this information to anticipate staffing needs until the tickets sell out.
You provide the manager with a quadratic equation that models the expected number of ticket sales for each day x.
(x = 1 is the day tickets go on sale).
Tickets = -0.6x2 + 12x + 11
Does the graph of the equation above open up or down?
Opens down
.
How did you determine whether the graph of the equation above opened up or down?
By looking at the coeeficient associated with the x^2 term.
positive:opens up
negative:opens down
.
Describe what happens to the ticket sales as time passes.
rises
reaches peak
then starts to decline
.
Use the quadratic formula to determine the last day that tickets will be sold.
Hint: Write your answer in terms of the number of days after ticket sales begin and you can’t have partial days.
21 days
.
Will tickets peak or be at a low during the middle of the sale?
Will peak... because it is a "quadratic" (parabola)
.
How do you know?
because it is a parabola that opens downward
.
After how many days will the peak or low occur?
Hint: Use the axis of symmetry formula.
x = -b/(2a)
x = -12/(2(-0.6))
x = -12/(-1.2)
x = 10 days
.
How many tickets will be sold on the day when the peak or low occurs?
Hint: Substitute your value from question 7 into the original quadratic equation.
Tickets = -0.6x2 + 12x + 11
Tickets = -0.6(10)^2 + 12(10) + 11
Tickets = -0.6(100) + 120 + 11
Tickets = -60 + 120 + 11
Tickets = 71
.
What are the coordinates of the vertex?
(10, 71)
How did you determine this?
How many solutions are there to the original quadratic equation given in question 3?
How do you know?